JustSaiyan:

if y=-2/5x-1/5 and y=3/2x-27/2, then what is solution? (7, 3) (-7, 3) (7, -3) (-7, -3)

Hanna:

(7, -3)

MARC:

we do not like fractions when solving simultaneous eqn... \(y=-\frac{2}{5}x-\frac{1}{5}\)---> multiply the eqn by 5 \(y=-\frac{3}{2}x-\frac{27}{2}\)--->multiply the eqn by 2 \(5y=-2x-1\) \(2y=-3x-27\) \(5y+2x=-1\)-->First eqn \(2y+3x=-27\)--> Second eqn Now,i'm using elimination method... Let's say we want to eliminate y,we must make both coefficient of the eqn the same... \(5y+2x=-1\)--> multiply the eqn by 2 \(2y+3x=-27\)-->multiply the eqn by 5 Both the eqns will look like this... \(10y+4x=-2\) \(10y+15x=-27\)

MARC:

\(\color{red}{10y}+4x=-2\) \(\color{red}{10y}-15x=-135\)

MARC:

since,both hv the same coefficient in each eqn... now,it is easier to eliminate y...

MARC:



MARC:

\(19x=133\) divide both sides by 19

MARC:

u will get \(x=7\)

MARC:

substitute x=7 into the first eqn.

MARC:

\(5y+2x=-1\)

MARC:

\(5y+2(7)=-1\) \(5y+14=-1\)

MARC:

subtract both sides by 14.

MARC:

\(5y=-14-1\) \(5y=-15\)

MARC:

Lastly,divide both sides by 5

MARC:

\(x=\frac{-15}{5}\)

MARC:

simplify it... \(x=\frac{-3\times5}{1\times5}\) \(x=\frac{-3}{1}\times\frac{5}{5}\) \(x=?\)

JustSaiyan:

That's a lot of math that looks suspiciously like a calculator, but does it answer my multiple choice question?

MARC:

Yes,it does... @JustSaiyan

Hanna:

yeah its (7, -3)

MARC:

yep,@Hanna is right! ^

JustSaiyan:

That was a lot of work. you could have just said that.

JustSaiyan:

But I get it. You wanna get your smart score up.

MARC:

i wanna u to understand better... ^ if i wanna get my smart score up,i could be at 100 already by just asking more question n spamming... xD

MARC:

want u*

JustSaiyan:

Lol thats how I got here. I asked history questions, lol